# MauveSim: the instrument simulator software for the Mauve missionMauveSim: the instrument simulator software for the Mauve mission

### v2

@@ Abstract
We present MauveSim, the instrument simulator software for Mauve, the latest mission from Blue Skies Space dedicated to time-domain stellar astronomy. The tool is designed to generate simulated stellar spectra, enabling the assessment of various scientific objectives, as well as determining limiting magnitudes and conducting signal-to-noise (S/N) analyses. MauveSim functions as an end-to-end simulator that takes an input stellar spectrum–either observed or synthetic–and produces a simulated observation based on the instrument’s performance and characteristics. The results of MauveSim have been validated against instrument performance data from extensive ground testing campaigns, ensuring that the software reflects the most up-to-date understanding of the payload performance. Accessible to all scientists involved in the mission, MauveSim serves as a crucial tool for target selection and observation planning.
@@ 1 Introduction
In this landscape, Blue Skies Space Ltd. (BSSL) has conceived Mauve, a UV mission designed to address the current gap in UV data availability. Mauve is a smallsat dedicated to time-domain spectrophotometry in the UV and Visible (UV-Vis) spectral range. Throughout its mission lifetime, Mauve will obtain time series of low-resolution spectra at $\mathscr{R}$ = 20-65 over a broad spectral range (200-700 nm), enabling the investigation of phenomena for which so far only fragmentary data have been collected.
@@
In this paper, we present MauveSim, the instrument simulator developed by BSSL for the Mauve mission, incorporating information from design choices and instrument performance evaluated during multiple ground testing campaigns. An instrument simulator is a software tool that allows users to simulate how an astronomical instrument will see a given astronomical source and thus allows scientists to propose observations in an informed way. The instrument simulator encodes the best understanding of the instrument’s characteristics and performance. As such, it will naturally develop over time, as the understanding of the instrument evolves, and it will be updated in flight when the final calibration will be achieved. The tool is available to all scientists participating in the Mauve mission. We offer interested users outside of the mission the opportunity to access the software upon request.
@@ 2 Payload components
The Mauve payload consists of three main components: a telescope, an optical fibre and a spectrometer. The payload has been designed with redundancy in mind, so the current baseline includes the provision of two fibre+spectrometer units to cope with a potential failure of the primary unit. The telescope and optical components are housed separately from the spacecraft platform equipment to minimise thermal variations throughout the orbit and reduce thermo-elastic distortions. The design, being developed by C3S and ISISpace, leverages heritage components where feasible and utilises the product lines and supply chains of both companies (Stotesbury et al., in prep.).
  • To mount the fibre between the telescope and the spectrometer, MediaLario modified the mount of the telescope in order for the fibre to placed at the telescope focus. The fibre will exit from the back of the primary mirror and connect to the spectrometer, aligning its output with the spectrometer aperture.
Table 1: Mauve’s spectrometer and detector specifications.
@@
Array Size 2048$\times$1 pix
Pixel Size 14$\times$200 $\mu$m
@@ 3 Instrument performance
Table 2: Instrument performance specifications.
@@
System’s optical efficiency 66%
To evaluate Mauve’s scientific performance, we used real measurements of the telescope, fibre, and spectrometer obtained during multiple test campaigns at MediaLario. Performance was measured in the visible spectrum (using a 636 nm filter) and in the UV wavelengths (using a 248 nm filter). The payload performance is primarily influenced by two key factors: the light throughput across the system and the detector’s dark current, which increases exponentially with temperature.
Regarding the first point, recent integrated payload tests revealed an overall optical efficiency of 66% throughout the system.
As for the second point, the thermal model provided by the suppliers assumes a detector operating temperature of $21^{\circ}\mathrm{C}$. While dedicated spacecraft instruments typically operate at cooler temperatures, Mauve’s COTS spectrograph is subject to higher temperatures due to design constraints. Ongoing work with the manufacturing partners aims to lower the detector temperature to a target of $13^{\circ}\mathrm{C}$. A refined thermal analysis, including a margin approach, suggests that this could halve the dark current and significantly improve instrument performance.
The example science simulations in Sec. 5 are based on conservative assumptions regarding the detector temperature and the system’s optical efficiency measured in the lab as presented in Table 2.
A third factor influencing and potentially limiting scientific performance is relying on payload data for pointing, rather than solely using star tracker information. This approach will be essential to compensate for the expected thermal distortion caused by the spacecraft crossing the Earth’s terminator during each orbit. Our concept consists in determining the fraction of time that the source’s light is actually collected by the fibre, as opposed to being lost when the source falls outside the fibre. The approach taken in defining the diameter of the fibre and thus the Field of View (FoV) has been to use the high-frequency pointing performance of the spacecraft. This has been modelled by ISISpace using the Moments of Inertia (MoI) calculations of the mature mechanical model and the ADCS controller and actuators in the system. The analysis has produced a Power Spectral Density (PSD) which predicts a high-frequency pointing stability of approximately 12” (10 s rms). We then used this information to conduct a time-domain analysis of the jitter, which helped us determine the fibre width. We selected a fibre diameter of 230 $\mu$m, corresponding to a half-cone FoV of approximately 47”. The algorithm developed by BSSL to correct for low-frequency pointing distortions is called “Payload In The Loop” (PITL) and it will be described in Bradley et al. (in prep.). However, the instrument simulator does not rely on the performance of the PITL and the results shown here assume that the spacecraft can lock onto and follow the target source selected for the simulation.
@@ 4 Workflow structure
MauveSim produces synthetic Mauve observations by applying a series of steps. These can be grouped in 4 broad categories.
@@ 4.1 Preparing the input spectrum
The simulator takes as input a source spectrum in the Mauve wavelength range (200-700 nm) expressed as a list of wavelengths paired with corresponding fluxes. The input spectrum can be either synthetic or the result of observations with other instruments. Stellar model SEDs, e.g. ATLAS9 models from Castelli & Kurucz [^1] https://wwwuser.oats.inaf.it/fiorella.castelli/ can be used to simulate Mauve observations, as long as they are provided in the format requested by the software. Alternatively, one can employ databases of observed stellar spectra. An example is the HST/STIS Low-resolution Stellar Library [^2] http://astro.wsu.edu/hststarlib/ (Pal et al., 2023) which provides more than 500 calibrated stellar spectra across the 200-1000 nm wavelength region at $\mathscr{R}$ $\sim$ 1000. Spectra from the STIS library are given in units of Å for the wavelength and $\text{erg}\,\text{s}^{-1}\,\text{cm}^{-2}\,\text{\AA}^{-1}$  for the flux. The flux is provided with an uncertainty as well.
Figure 1: Magnitude comparison for the star HD 143459.
Once the user selects the input spectrum, this is downsampled if it contains more than 2048 data points. This interpolation is necessary to ensure that each wavelength data point is consistent with the association of the detector array with the range created by the combination of the grating plus order sorting filter.
The star taken as an example here, HD 143459, has a V magnitude of 5.526. To simulate the same spectrum but with a different magnitude, the flux needs to be scaled according to
@@
where $F_{0}$ is the original flux and $\Delta m$ = $m_{\text{final}}$ $m_{\text{initial}}$, i.e. the difference between the final wanted magnitude (e.g. $V_{\text{mag}}$ = 3) and the initial magnitude (in this case $V_{\text{mag}}$ = 5.526). Fig. 1 displays the original spectrum of HD 143459 retrieved from the STIS stellar database, alongside its version scaled to a visual magnitude of 3.
Figure 2: Zodiacal sky background ($V_{\text{mag}}$ $\text{arcsec}^{-2}$) as a function of helio-ecliptic coordinates. Table from COS Instrument Handbook.
@@
To calculate the zodiacal noise, we use a table of approximate zodiacal sky background as a function of helio-ecliptic coordinates. First, the user is prompted to provide the RA, Dec and time of observation (UTC) to check if the target is within the observing zone or in the solar-avoidance zone. The equatorial coordinates RA and Dec are first converted to geocentric ecliptic coordinates using astropy modules. In the geo-ecliptic coordinate system, the zero longitude is defined by the vernal equinox. However, the zodiacal light is a function of the Sun’s position. This is what defines the zero longitude in the helio-ecliptic coordinate system. Thus, by using the input time of observation, the Sun’s position from Earth’s perspective is calculated at the given time. Then, the Sun’s position is transformed to geocentric ecliptic coordinates (Leinert et al., 1998). The geocentric ecliptic coordinates are converted to helio-ecliptic coordinates according to:
@@
where $\lambda$ represents longitude, $\beta$ represents latitude and $\lambda_{\odot}$ represents the Sun geo-ecliptic longitude. Therefore, the level of zodiacal light is a function of the position of the star in the sky and of the time it is observed (more or less close to the Sun). The helio-ecliptic coordinates thus defined are then assigned the corresponding level of zodiacal background, using the reference table in Fig. 2.
The zodiacal sky background is given in units of $V_{\text{mag}}$ $\text{arcsec}^{-2}$. To convert it into flux, we follow the steps below. Assuming that the zodiacal background has a signature similar to the Solar spectrum, we take the spectrum of HD 126614 (a G8 star) and scale its magnitude to
@@
$V_{\text{total}}$ is calculated by taking into account the magnitude of the zodiacal light read off from the table ($V_{\text{zodi}}$) and the field of view of the telescope $(A_{\text{FoV}})$, which is 47.46 arcsec in diameter, 1769 arcsec2 in area. Equation 3 comes from the relationship between surface brightness (magnitude per square arcsecond) and the total magnitude over a given area. The surface brightness ($V_{\text{zodi}}$) is the apparent magnitude per square arcsecond. It describes how bright an object appears spread over one square arcsecond of the sky. The total magnitude ($V_{\text{total}}$) is the total apparent magnitude of an object when considering the entire area it covers. When a surface brightness is given in square arcseconds, it is necessary to account for how the flux adds up over that area. The total flux is the surface brightness flux multiplied by the area:
@@
If the spectrum has a coverage beyond 200-700 nm, the software considers only the wavelengths between 200 and 700 nm.
Fig. 3 illustrates how the target spectrum is altered by the inclusion of the zodiacal background. While the two spectra appear visually indistinguishable, the residuals reveal that the difference corresponds to a solar-type spectrum with an average $V_{\text{mag}}=22$.
@@ 4.2 Calculating the source signal
Figure 4: The spectrum of HD 143459, obtained using HST/STIS, has been convolved to Mauve’s resolution using a combination of Gaussian and top-hat kernels. The graph illustrates a resolution comparison. However, it is important to note that the spectrum in green does not represent the flux levels achievable by Mauve.
To express the spectrum according to the resolution of the instrument, MauveSim convolves the result of step 4.1 with a kernel that downgrades the data to the resolution achievable by Mauve. Here we convolve the stellar spectrum with a kernel (a smoothing function), that is defined by the instrument line function (ILF).
@@
[…]
@@
The instrument line function, determines how the original spectral data gets broadened. The Mauve ILF features two components: a “top hat” element due to the fibre and a Gaussian component due to the spectrograph. A one-dimensional top hat kernel centred at $\lambda=0$ with width $2a$ (i.e., it extends from $-a$ to $a$) is defined as:
@@
According to the spectrometer specifications, for a grating of 600 lines/mm and a 500 $\mu$m slit, the resolution is 108 Å. This is the width of the boxcar kernel.
@@
$k(\lambda)=\frac{1}{\sigma\sqrt{2\pi}}e^{-\frac{\lambda^{2}}{2\sigma^{2}}}\ .$ (11)
The standard deviation $\sigma$ determines the width of the kernel, which directly influences the degree of resolution reduction. A broader kernel, corresponding to a larger $\sigma$, results in a more pronounced smoothing effect, leading to a greater reduction in resolution.
Mauve has constant resolution with wavelength, meaning that the value of $\Delta\lambda$, the smallest resolvable wavelength difference, is the same throughout the entire wavelength range of the instrument. For Mauve, $\Delta\lambda$approx 105 Å, which determines the FWHM of the Gaussian kernel. The top hat kernel and Gaussian kernel are first convolved together, then their convolution is applied to the data. The convolution order does not matter as it is a commutative operation. Fig. 4 displays the resolution achievable by Mauve.

### v3

# MauveSim: the instrument simulator software for the Blue Skies Space Mauve satelliteMauveSim: the instrument simulator software for the Blue Skies Space Mauve satellite

@@ Abstract
We present MauveSim, the instrument simulator software for Mauve, the latest mission from Blue Skies Space dedicated to time-domain stellar astronomy. MauveSim functions as an end-to-end simulator, employing the most up-to-date knowledge of the instrument’s performance and characteristics that will be reviewed and updated after commissioning. The software accepts a stellar spectrum–either observed or synthetic–as input and produces a simulated observation. The tool thus enables the assessment of various scientific objectives, as well as determining limiting magnitudes and conducting signal-to-noise (S/N) analyses. The results of MauveSim have been validated against instrument performance data from extensive ground testing campaigns, ensuring that the software reflects the most up-to-date understanding of the payload performance. Accessible to all scientists involved in the mission, MauveSim serves as a crucial tool for target selection and observation planning.
@@ 1 Introduction
In this landscape, Blue Skies Space Ltd. (BSSL) has conceived Mauve, a 16U smallsat designed to address the current gap in UV data availability. The mission is dedicated to time-domain spectrophotometry in the UV and Visible (UV-Vis) spectral range. Throughout its lifetime, Mauve will obtain time series of low-resolution spectra at $\mathscr{R}$ = 20-65 over a broad spectral range (200-700 nm), enabling the investigation of astrophysical phenomena for which so far only fragmentary data have been collected. Even a small satellite like Mauve will be able to provide much needed coverage in the UV and optical bands, which can complement existing spectroscopic stellar data at longer wavelengths. Mauve can effectively survey stellar targets in the UV and visible, enabling detailed characterisation of stellar activity and variability, key for interpreting exoplanet spectroscopic observations. The capabilities of Mauve in monitoring stellar targets are especially valuable in the current space observatories landscape, given the James Webb Space Telescope’s lack of sensitivity below 600 nm (Greene et al., 2016). Additionally, Mauve will be capable of obtaining near-ultraviolet spectra and constrain the variability of stars identified as high-priority targets for upcoming exoplanet-dedicated missions such as Ariel (Tinetti et al., 2018, 2021) and Twinkle (Edwards et al., 2019; Stotesbury et al., 2022) and the future Habitable Worlds Observatory (Mamajek & Stapelfeldt, 2024).
@@
The spacecraft will be launched in a low Earth Sun-synchronous orbit at 500 km of altitude, with an orbital period of approximately 95 minutes. As a result, Mauve will intersect regularly with the South Atlantic Anomaly (SAA). BSSL’s intention is to use the commissioning phase to qualify the performance during SAA crossing and determine if there is an impact on the science measurements. If needed, BSSL will determine a flux density cut-off above which to cease payload operations and pause the scheduled observations.
In this paper, we present MauveSim, the instrument simulator developed by BSSL for the Mauve mission. Instrument simulators are essential tools for mission preparation. They are designed to replicate the performance of astronomical instruments, incorporating information from design choices and the best understanding of the instrument’s characteristics and performance evaluated during ground testing campaigns (Bernie Simon, 1986; Dobke et al., 2010; Oscar A. Gonzalez, 2019). As such, an instrument simulator naturally develops over time, as the understanding of the instrument evolves, and it is updated in flight when the final calibration is achieved. Simulators also enable scientists to test, propose in an informed way, and analyse observations at all stages–before, during, and after data acquisition–providing insights into the influence of instrumental effects on the data. Moreover, simulators are vital for optimising observation strategies and validating data reduction pipelines.
MauveSim is written entirely in Python and accessible on Stardrive, a web-based platform developed by BSSL that integrates simulation results, data management, and collaboration tools in a unified environment. Stardrive is available for all scientists participating in the Mauve mission. At BSSL, we are supporting MauveSim as an actively maintained tool within the Stardrive platform, with regular updates driven by mission requirements and feedback from the Mauve science consortium. The instrument simulator is available to Mauve users through a secure, web-based interface as part of our mission planning and data management suite. We also offer ad-hoc access upon request. The software is maintained under version control and the software version is indicated in the web-interface. Our long-term strategy additionally includes ongoing support and continuous improvements to ensure MauveSim remains reliable, scientifically accurate, and aligned with the evolving needs of the mission.
This paper is organised as follows. Section 2 presents the Mauve’s payload design and the characteristics of its components. More details on the simulator validation activities are available in Section 3, where we also describe the limitations we encountered during the testing campaign when trying to characterise the payload as accurately as possible. The description of the algorithm is presented in Section 4, which explains in detail the assumptions, input data, implementation and output of each step in the workflow. Finally, Section 5 presents the application of the instrument simulator to one of the Mauve "flagship" science cases, showing how the software can be used to understand limiting magnitudes, and thus perform science definition.
@@ 2 Payload components
Figure 1: Illustration of Mauve optical diagram (not to scale).
Table 1: Mauve’s spectrometer, detector and telescope specifications.
@@
Wavelength coverage 200-700 nm
Wavelength range per pixel ($\Delta x_{\lambda}$) 0.25-0.30 nm
[…]
@@
Telescope Parameters
Field of View 47.46” radius
Collecting area ($A$) 122 cm2
The Mauve payload consists of three main components: a telescope, an optical fibre and a spectrometer. The payload has been designed with redundancy in mind, so the current baseline includes the provision of two fibre+spectrometer units to cope with a potential failure of the primary unit. The telescope and optical components are housed separately from the spacecraft platform equipment to minimise thermal variations throughout the orbit and reduce thermo-elastic distortions. The design, developed by C3S and ISISPACE, leverages heritage components where feasible and utilises the product lines and supply chains of both companies (Stotesbury et al., in prep.).
  • To mount the fibre between the telescope and the spectrometer, MediaLario modified the mount of the telescope to allow the fibre to be placed at the telescope focus. The fibre will exit from the back of the primary mirror and connect to the spectrometer, aligning its output with the spectrometer aperture. The payload optical design is provided in Fig. 1.
@@ 3 Instrument performance
Table 2: Payload performance specifications.
@@
Light loss across the system 34%
[…]
We evaluated Mauve’s performance through a series of laboratory measurements of the telescope, fibre, and spectrometer, obtained during several testing campaigns at MediaLario. During the functional tests, the payload’s performance was assessed at multiple stages along the optical chain. Measurements were performed using two narrow-band filters: one in the ultraviolet region at 248 nm and another in the visible wavelengths at 636 nm. To validate the results, measured signal outputs were compared against expected values based on the efficiency curves of each payload component, as provided by the manufacturers. A consistent signal level of approximately 34% below the expected signal value was observed across all tests, indicating a lower efficiency of the payload components. While we don’t expect this to impact results significantly, we note that the laboratory tests we performed had limitations in producing a perfectly collimated beam, potentially impacting the coupling to the fibre. Furthermore, the estimation of UV performance is less accurate than that of visible wavelength performance, due to wavelength dependence in scattering properties of the telescope mirrors compared to band integrated measurements of efficiency.
The conservative thermal model supplied by the manufacturers assumes a detector operating temperature of $21^{\circ}\mathrm{C}$. While dedicated spacecraft instruments typically operate at cooler temperatures, Mauve’s COTS spectrograph is subject to higher temperatures due to design constraints. Ongoing work with the manufacturing partners aims to lower the detector temperature to a target of $13^{\circ}\mathrm{C}$. A refined thermal analysis, including a margin approach, suggests that this could halve the dark current and significantly improve instrument performance.
The example science simulations in Sec. 5 are based on the conservative thermal model while the system’s light loss is the one measured in the lab. These parameters are presented in Table 2.
[…]
@@ 4 Workflow structure
MauveSim produces simulated Mauve observations by applying a series of steps. These can be grouped in 4 broad categories.
@@ 4.1 Preparing the input spectrum
In the MauveSim interface we provide an integrated library of stellar spectra that enables the users to select appropriate inputs for their simulations. These stellar spectra are sourced from the HST/STIS Stellar Spectral Library [^1] http://astro.wsu.edu/hststarlib/ (Pal et al., 2023) which provides more than 500 calibrated stellar spectra across the 170-1000 nm wavelength region at $\mathscr{R}$ $\sim$ 1000. Spectra from the STIS library are given in units of Å for the wavelength and $\text{erg}\,\text{s}^{-1}\,\text{cm}^{-2}\,\text{\AA }^{-1}$ for the spectral flux density.
Alternatively, users may upload their own spectrum via the simulator interface, provided the input data adheres to specified guidelines. The required input file must be in ASCII format, with a .txt or .dat extension. The file must contain two numerical columns separated by spaces, tabs or commas. Column 1 must contain the wavelength array in units of nm, while the second column must contain the spectral flux density array in units of $\text{erg}\,\text{s}^{-1}\,\text{cm}^{-2}\,\text{\AA }^{-1}$. The spectrum must span at least 180–750 nm, which is the detector’s operational spectral band. While the detector can capture light beyond the spectrometer’s wavelength range of 200-700, the payload’s efficiency outside of this range is generally too poor to produce scientifically useful data. To account for this, the software restricts the output to the 200–700 nm range.
Figure 3: Original spectrum of HD 37202 (Vmag=3.03) compared to its rescaled version at Vmag=5. This example demonstrates the instrument simulator’s capability to modify the input flux to a user’s specified magnitude.
Once the user selects the input spectrum, this is downsampled if it contains more than 2048 data points, the number of detector pixels in the spectral direction. This interpolation is necessary to ensure that each wavelength data point is consistent with the association of the detector array with the range created by the combination of the grating plus order sorting filter.
[…]
The software includes an option to simulate the input spectrum at a different magnitude. This is achieved by scaling the flux according to
@@
where $F_{0}$ is the original flux and $\Delta m$ = $m_{\text{final}}$ - $m_{\text{initial}}$, i.e. the difference between the final wanted magnitude (e.g. $V_{\text{mag}}$ = 5) and the initial magnitude (in this case $V_{\text{mag}}$ = 3.03). To illustrate the magnitude rescaling option, Fig. 3 displays the original spectrum of HD 37202 retrieved from the STIS stellar database, alongside its version scaled to a visual magnitude of 5.
Table 3: Zodiacal sky background ($V_{\text{mag}}$ $\text{arcsec}^{-2}$) as a function of helio-ecliptic coordinates. Table from the COS Instrument Handbook.
Helio-ecliptic Longitude Helio-ecliptic Latitude
0 15 30 45 60 75 90
180 22.1 22.4 22.7 23.0 23.2 23.4 23.3
165 22.3 22.5 22.8 23.0 23.2 23.4 23.3
150 22.4 22.6 22.9 23.1 23.3 23.4 23.3
135 22.4 22.6 22.9 23.2 23.3 23.4 23.3
120 22.4 22.6 22.9 23.2 23.3 23.3 23.3
105 22.2 22.5 22.9 23.1 23.3 23.3 23.3
90 22.0 22.3 22.7 23.0 23.2 23.3 23.3
75 21.7 22.2 22.6 22.9 23.1 23.2 23.3
60 21.3 21.9 22.4 22.7 23.0 23.2 23.3
45 SA SA 22.1 22.5 22.9 23.1 23.3
30 SA SA SA 22.3 22.7 23.1 23.3
15 SA SA SA SA 22.6 23.0 23.3
0 SA SA SA SA 22.6 23.0 23.3
@@
To calculate the zodiacal noise, we use a table of approximate zodiacal sky background as a function of helio-ecliptic coordinates. The user is prompted to provide the RA, Dec and time of observation (UTC) to check if the target is within the observing zone or in the solar-avoidance zone. The equatorial coordinates RA and Dec are first converted to geocentric ecliptic coordinates. In the geo-ecliptic coordinate system, the zero longitude is defined by the vernal equinox. However, the zodiacal light is a function of the Sun’s position. By using the input time of observation, the Sun’s position from Earth’s perspective is calculated at the given time and is transformed to geocentric ecliptic coordinates (Leinert et al., 1998). The geocentric ecliptic coordinates are converted to helio-ecliptic coordinates according to:
@@
where $\lambda$ represents longitude, $\beta$ represents latitude and $\lambda_{\odot}$ represents the Sun geo-ecliptic longitude. Therefore, the level of zodiacal light is a function of the position of the star in the sky and of the time it is observed (more or less close to the Sun). The helio-ecliptic coordinates thus defined are then assigned the corresponding level of zodiacal background, using the reference Table 3.
The zodiacal sky background is given in units of $V_{\text{mag}}$ $\text{arcsec}^{-2}$. To convert it into flux, we assume that the zodiacal background has a spectrum similar to the solar one, which is a reasonable assumption considering the wavelength coverage of Mauve (Kelsall et al., 1998; Pascale et al., 2015). On MauveSim we use the spectrum of HD 87646 (a G2V star) and scale its magnitude to
@@
$V_{\text{total}}$ is calculated by taking into account the magnitude of the zodiacal light read off from the table ($V_{\text{zodi}}$) and the field of view of the telescope $(A_{\text{FoV}})$, 47.46 arcsec in radius, corresponding to an area of 7072 arcsec2 $(A_{\text{FoV}})$. Equation 3 comes from the relationship between surface brightness (magnitude per square arcsecond) and the total magnitude over a given area. The surface brightness ($V_{\text{zodi}}$) is the apparent magnitude per square arcsecond. It describes how bright an object appears spread over one square arcsecond of the sky. The total magnitude ($V_{\text{total}}$) is the total apparent magnitude of an object when considering the entire area it covers. When a surface brightness is given in square arcseconds, it is necessary to account for how the flux adds up over that area. The total flux is the surface brightness flux multiplied by the area:
@@
Fig. 4 illustrates how the target spectrum is altered by the inclusion of the zodiacal background. While the two spectra appear visually indistinguishable, the residuals reveal that their difference corresponds to a solar-type spectrum with an average $V_{\text{mag}}=22$.
@@ 4.2 Calculating the source signal
Figure 5: The spectrum of HD 37202 observed with HST/STIS is shown in blue. In green the same spectrum has been convolved to Mauve’s instrument line function. The spectrum in green does not represent the flux levels achievable by Mauve.
To express the spectrum according to the resolution of the instrument, MauveSim convolves the result of step 4.1 with a kernel that downgrades the data to the resolution of Mauve. Here we convolve the stellar spectrum with a kernel representing the instrument line function (ILF).
@@
$y(\lambda)=(x*k)(\lambda)=\int_{0}^{\infty}x(\lambda^{\prime})k(\lambda-\lambda^{\prime})\,d\lambda^{\prime}\ .$ (8)
@@
The instrument line function, determines how the original spectral data gets broadened. The Mauve ILF features two components: a “top hat” element due to the fibre and a Gaussian component due to the spectrograph. A one-dimensional top hat kernel centred at $\lambda=0$ with width $2a$ (i.e., it extends from a-a to $a$) is defined as:
@@
$k(\lambda)=\frac{1}{\sigma\sqrt{2\pi}}e^{-\frac{\lambda^{2}}{2\sigma^{2}}}\ ,$ (11)
where the standard deviation $\sigma$ determines the width of the kernel.
Mauve has constant resolution with wavelength, meaning that the value of $\Delta\lambda$, the smallest resolvable wavelength difference, is the same throughout the entire wavelength range of the instrument. For Mauve, $\Delta\lambda$ $\approx$ 105 Å, which determines the FWHM of the Gaussian kernel. The top hat kernel and Gaussian kernel are first convolved together, then their convolution is applied to the data. The convolution order does not matter as it is a commutative operation. Fig. 5 displays the spectrum of HD 37202 after convolution with the Mauve instrument line function.